OpenStudy (anonymous):

solve (x-3)^2 - 4 = 0. 4a) solve by taking square roots

OpenStudy (anonymous):

i am not sure but it looks like (x-3)^2-4=0 (x-3)^=4 x-3=2 x=5

OpenStudy (anonymous):

oh okay thank you very much